PHOSPHORUS CHEMISTRY IN CIRCUMSTELLAR ENVELOPES: PN IN IRC+10216, VY CMA, AND CRL2688

\underline{ALDO J. APPONI}, STEFANIE N. MILAM, DEWAYNE T. HALFEN, LUCY M. ZIURYS, UNI-VERSITY OF ARIZONA, STEWARD OBSERVATORY, ARIZONA RADIO OBSERVATORY, 933 N CHERRY AVENUE, TUCSON, AZ 85721.

Currently the role of phosphorus in circumstellar chemistry is uncertain. The only phosphorus bearing species detected to date is CP, which has been identified only in IRC+10216. Recent observations using the facilities of the Arizona Radio Observatory (ARO) have suggested that PN is a relevant circumstellar molecule as well. Using the ARO 12 meter and SMT telescopes to observe the J=2-1, 3-2, and 5-4 transitions, PN has now been identified in IRC+10216, VY CMa, and tentatively, CRL2688. Typical abundances in these sources are 1 − 4 × 10^{−8}, relative to H_{2}. Observations towards additional circumstellar envelopes are currently being conducted, including CIT6, CRL618, TX Cam and NGC7027.